# Towards Automated Semantic Interpretability in Reinforcement Learning via Vision-Language Models

### v1

@@ III-B Tree-based RL Training
After obtaining the Extractor CNN, we train an interpretable tree-based model with RL to connect the semantic features with action outputs. During RL training, our framework utilizes the pre-trained Extractor CNN to extract semantic features from image observations, as shown in Fig. 1. The policy network takes these semantic features as input and predicts a distribution over actions, from which an action is sampled and applied to the controlled agent at each time step. Unlike traditional RL algorithms that use black-box models as the policy network, we use an Interpretable Control Tree (ICT) as our policy network to achieve transparent decision-making and human-aligned reasoning while remaining differentiable for end-to-end optimization.
@@
Interpretable Control Tree: Interpretable Control Tree (ICT) is adapted from Interpretable Continuous Control Tree (ICCT) proposed by Paleja et al.  [6]. The ICT is structured similarly as ICCT, where each decision node determines the routing of input features, $\mathbf{x}$, to sparse linear controllers at the leaf nodes.
@@
Each decision node, $i$, is parameterized by weights $\mathbf{w}_{i}$, a bias $b_{i}$, and a steepness parameter $\alpha$ to output a fuzzy Boolean in $[0,1]$, as shown in Eq. 3.
@@
Unlike ICCT, ICT naturally handles continuous and discrete action space. In the discrete case, leaf nodes in ICT output the possibility of taking each action as a Probability Mass Function (PMF). If there are six possible actions in the discrete action space, then there will be six sparse linear functions of the input features followed by a softmax function at each leaf node. And for the continuous action space, leaf nodes output the Probability Density Function (PDF) over possible actions.
IV Experiments
IV-A Environment
We evaluate our approach on two Atari environments: Skiing and Boxing, and a classic control domain: Cart-pole, where we use image snapshots instead of the built-in feature vector as the observation space. The OpenAI Atari platform is preferable because all their environments provide raw image state spaces without any pre-defined feature vectors or feature sets. This ensures that the VLM engages in genuine reasoning over the environment to identify key semantic features, rather than relying on prior internet-based knowledge, a phenomenon referred to as training data memorization [38].
IV-B Feature Extraction
IV-B1 VLM Deducing Key Semantic Features.
@@ IV-B Feature Extraction
For this step, we utilize GPT-4o, the current state-of-the-art in language reasoning. Across all three environments, the VLM consistently produces a deterministic set of semantic features for decision-making, achieving performance comparable to human reasoning. For all three environments, we query the VLM 50 times and use majority vote to get the most prominent answer from the VLM. We present the identified feature set for the Skiing environment in Fig. 1. Other domains are included in Appendix VII-D.
(a) Boxing
(b) Cartpole
(c) Skiing
Figure 2: Results across three domains. Higher scores are better.
IV-B2 VLMs for Feature Extraction Code
@@
For this step, we utilize claude-3.5-sonnet-20241022, the current state-of-the-art in code generation. To address common challenges with VLM-generated code, such as non-functionality or runtime errors, we query the VLM 50 times using the same prompt (temperature = 0 for reproducibility) and select the most reliable feature extraction code through majority voting.
@@
The generated codes are evaluated on a randomly collected dataset of 100 images and filtered based on the following criteria: (1) the code must compile and execute successfully and (2) it must produce diverse feature outputs rather than constant or repetitive values across the dataset. Candidates producing identical outputs are grouped into clusters, and the largest cluster is chosen as the final feature extraction code. Table I presents statistics from this process, including code failure rates, exclusions due to constant outputs, and majority selection coverage. The majority selection coverage appears low because features related to relative distances are normalized differently, while features related to object positions remain identical.
It is important to note that while Step 1, which generates $\mathcal{G}_{s}$ for feature extraction from a single image frame, is essential for all tasks, Step 2, which generates $\mathcal{G}_{m}$ for feature extraction across multiple image frames, is only necessary when the task’s features depend on temporal information. In the Boxing environment, the feature space does not depend on temporal inference, making Step 2 unnecessary. Consequently, Step 2 is marked as empty for Boxing in Table I. However, the VLM was still queried 50 times, and in all 50 cases, it successfully determined that no code generation was needed, achieving a 100% success rate under this metric.
IV-B3 Extractor CNN Performance Analysis
@@
In Step 3 of our methodology, we address the inefficiencies of the VLM-generated feature extraction code for single-image frames, $\mathcal{G}_{s}$, by training an Extractor CNN as its replacement. While $\mathcal{G}_{s}$ can be executed directly on the dataset, the processing pipeline is inherently non-batchable, as it handles image segmentations independently (see Appendix VII-B for examples). This non-batchability severely impacts RL training, causing slowdowns of 200–2000× compared to batchable operations (Extractor CNN), as demonstrated in Table III. From practical experience, this would translate to training times exceeding multiple weeks on an H40-level GPU for a single policy to converge, making this infeasible.
@@
To constrain computing time to a reasonable range, we train an Extractor CNN to replace $\mathcal{G}_{s}$ and efficiently compute $\mathcal{F}_{s}$, the features derived from a single image frame. The Extractor CNN is specifically designed to resolve the slowdown caused by segmentation-based analysis in $\mathcal{G}_{s}$. For each environment, we train an Extractor CNN on 7,000 images collected from diverse trajectories, validate it on 2,000 images, and test it on 1,000 images. The models achieved a mean squared error (MSE) consistently below 0.02 for each feature in the output vector when compared to the VLM-generated ground truth. Table II provides detailed performance statistics. These results confirm that the Extractor CNN is a reliable and computationally efficient replacement for $\mathcal{G}_{s}$.
TABLE III: Runtime speedup for processing a batch of images across different environments on the same device (averaged over 10 iterations). Results denoted with * are derived from a different device with a better gpu.
Env. Batch VLM (s) CNN (s) Speedup
Skiing 8 0.870 0.011 81.02x
16 1.745 0.003 523.37x
32 3.523 0.005 743.77x
64 7.055 0.008 877.90x
128 14.449 0.021 693.74x
Cart-pole 8 12.726 0.017 738.57x
16 25.679 0.040 637.06x
32 50.204 0.066 763.96x
64 99.660 0.111 898.09x
128* 254.484 0.097 2617.22x
Boxing 8 0.400 0.017 23.07x
16 0.785 0.004 191.03x
32 1.581 0.007 238.51x
64 3.146 0.009 344.34x
128 6.240 0.025 254.40x
IV-C Training Results
Figure 3: Visualization of ICT in the Skiing Environment.
Figure 4: Demonstration of the discrepancy between the physics simulation and the visualization.
@@ IV-C Training Results
  • VLM-DT Heuristic: A baseline where VLM generates a heuristic tree to connect the semantic feature space with the action space.
@@
Across all three environments, VLM-ICT consistently outperforms other uninterpretable models (MLP) built on the semantic feature space, and demonstrates performance close to the CNN+RL black-box benchmark representing the upper limit of task performance. Notably, in the Cart-pole environment, VLM-ICT even outperforms the black-box CNN model. This is because the CNN fails to accurately capture the pole’s angle, a critical feature for formulating an effective policy. These consistent results highlight the robustness and effectiveness of our proposed method.
The learned ICT in Skiing domain is shown in Fig. 3. This tree was generated fully autonomously using the VLM $\rightarrow$ CNN $\rightarrow$ RL pipeline. The learned trees (ICT) are interpretable both subjectively and objectively, provided they remain small enough, as demonstrated in a previous user study [6]. For the visualization of our learned ICT in Boxing and Cart-pole domains, please refer to Appendix VII-A.
V Limitations and Future Works
In the Cart-pole environment, we observe that the performance of trained VLM-ICT and VLM-MLP has yet to match that of RL policies trained using ground truth features [39, 6]. This discrepancy stems from the gap between the underlying physics simulation and its visual representation, as illustrated in Fig. 4. While the trained Extractor CNN’s predictions closely approximate the feature vectors extracted by the VLM-generated code, they still exhibit deviations from the ground truth feature vectors. This highlights a fundamental challenge in semantic RL from visual observations, where the extracted features may not perfectly align with the underlying physical state. Despite this limitation, our framework still outperforms the CNN-based policy, demonstrating the promise of VLMs in RL. To the best of our knowledge, we are the first to attempt solving the Cart-pole problem using image-based observations, demonstrating a novel and promising direction for interpretable RL.
@@ V Limitations and Future Works
Future works could explore RL-based backpropagation to fine-tune the VLM models, enabling large language models to iteratively refine their outputs by learning from feedback generated through interactions with RL agents [40]. This approach has the potential to enable VLMs to identify discrepancies between the underlying physics simulation and its visual representation, leading to improved feature extraction for downstream RL tasks.
VI Conclusion

### v2

@@ 3.2 Tree-based RL training
After obtaining the Extractor Module, we train an interpretable tree-based model with RL to connect the semantic features with action outputs. During RL training, our framework utilizes the pre-trained Extractor Module to extract semantic features from image observations, as shown in Fig. 1. The policy network takes these semantic features as input and predicts a distribution over actions, from which an action is sampled and applied to the controlled agent at each time step. Unlike traditional RL algorithms that use black-box models as the policy network, we use an Interpretable Control Tree (ICT) as our policy network to achieve transparent decision-making and human-aligned reasoning while remaining differentiable for end-to-end optimization.
3.2.1 Interpretable control tree
@@
Interpretable Control Tree (ICT) is adapted from Interpretable Continuous Control Tree (ICCT) proposed by Paleja et al.  [18]. The ICT is structured similarly as ICCT, where each decision node determines the routing of input features, $\mathbf{x}$, to sparse linear controllers at the leaf nodes.
@@
Each decision node, $i$, is parameterized by weights $\mathbf{w}_{i}$, a bias $b_{i}$, and a steepness parameter $\alpha$ to output a fuzzy Boolean in $[0,1]$, as defined in Eq. 3.
@@
Unlike ICCT, ICT supports both discrete and continuous actions without re-architecture, which helps generalizability. In the discrete case, leaf nodes in ICT output the possibility of taking each action as a Probability Mass Function (PMF). If there are six possible actions in the discrete action space, then there will be six sparse linear functions of the input features followed by a softmax function at each leaf node. And for the continuous action space, leaf nodes output the Probability Density Function (PDF) over possible actions.
4 Experiments
4.1 Environment
We evaluate our approach across three diverse environments: Atari Skiing, a fast-paced pixel-art game requiring gate navigation; BabyAI-GoToRedBall, a symbolic grid world for instruction following; and OpenAI Gym Highway, a 2D environment for high-level autonomous driving decisions like lane switching and speed control. The Atari Skiing environment provides raw image state spaces without any pre-defined feature vectors or feature sets. BabyAI and Highway offer default feature vectors, but these vectors describe the full state of the environment rather than focusing on the task-relevant semantic features needed for decision making. By selecting these environments, we ensure that the VLM engages in genuine reasoning over the environment to identify key semantic features, rather than relying on prior internet-based knowledge, a phenomenon referred to as training data memorization [5].
4.2 Feature extraction
4.2.1 VLM deducing key semantic features.
@@ 4.2 Feature extraction
For this step, we utilize ChatGPT-4o-latest, one of the current leading models in language reasoning. Across all three environments, the VLM consistently produces a deterministic set of semantic features for decision-making, achieving performance comparable to human reasoning. For all three environments, we query the VLM 50 times and use majority vote to get the most prominent answer from the VLM. A complete set of those features are included in Appendix A.2.2.
4.2.2 VLM generating Feature Extraction code.
@@
For this step, we utilize claude-3.5-sonnet-20241022, the current leading model in code generation. To address common challenges with VLM-generated code, such as non-functionality or runtime errors, we query the VLM 50 times using the same prompt (temperature = 0 for reproducibility) and select the most reliable feature extraction code through majority voting.
@@
The generated codes are evaluated on a randomly collected dataset of 100 images and filtered based on the following criteria: (1) the code must compile and execute successfully and (2) it must produce diverse feature outputs rather than constant or repetitive values across the dataset. Candidates producing identical outputs are grouped into clusters, and the largest cluster is chosen as the final feature extraction code. Details are included in Appendix A.2.
4.2.3 Extractor module performance analysis
@@
In Step 3 of our methodology, we address the inefficiencies of the VLM-generated feature extraction code for single-image frames, $\mathcal{G}_{s}$, by training an lightweight Extractor Module as its replacement. While $\mathcal{G}_{s}$ can be executed directly on the dataset, the processing pipeline is inherently non-batchable, as it handles image segmentations independently. This non-batchability severely impacts RL training, causing slowdowns of around 20000× compared to batchable operations (our CNN Extractor Module), as demonstrated in Table 3. From practical experience, this would translate to training times exceeding multiple weeks on an H40-level GPU for a single policy to converge.
@@
We constrain computing time to a reasonable range with a lightweight Extractor Module, resolving the slowdown caused by segmentation-based analysis in $\mathcal{G}_{s}$. For each environment, we train an CNN on 7,000 images collected from diverse trajectories, validate it on 2,000 images, and test it on 1,000 images. Compared to the VLM-generated ground truth, our models achieve high accuracy (typically >0.95) on most binary features, while MSE remains low (<0.005) for the majority of continuous features, with one exception where error is notably higher. Table 1 provides detailed performance statistics. These results confirm that a lightweight vision model can serve as a reliable and computationally efficient replication for VLM generated extraction code $\mathcal{G}_{s}$.
Table 1: Performance of the Extractor Module across environments, compared with ground truth from VLM-generated code. Continuous components are evaluated with MSE; boolean components are evaluated with accuracy. Dashes indicate vacant entries due to varying feature lengths and types across environments. Full feature descriptions are provided in Appendix A.2.2.
Environment Feature Vector Indices (1–7)
1 2 3 4 5 6 7
Skiing (MSE) 0.00023 0.00025 0.0020 0.0039 0.0024 0.00034
BabyAI (MSE) 0.0001 0.0001
BabyAI (Acc) 1.000 0.979 0.975 0.972
Highway (MSE) 0.0000 0.0000 1.153 0.0268 0.0209
Highway (Acc) 0.988 0.948
4.3 Training results
@@
(a) Atari Skiing
(b) BabyAI: GoToRedBall
(c) OpenAI Gym Highway
Table 2: Runtime speedup on Skiing.
Other domains in Appendix A.3.
Batch VLM Code CNN Speedup
8 19.056 0.001 13831.3x
16 38.288 0.002 17325.6x
32 76.212 0.004 19935.0x
64 151.836 0.007 21692.9x
128 308.468 0.014 22772.4x
Figure 3: Results across three domains. Higher scores are better. Runtime comparison shows speedup of CNN-based extractor over VLM code.
In addition to standard learning-based baselines, we also explore whether a Vision-Language Model (VLM) can directly perform action selection via prompting, given access to a semantic feature space it previously constructed. We prompt the VLM to synthesize an if-else style decision tree and obtain two variants of this baseline:
  • VLM-DT (ChatGPT-4o-latest): A baseline where ChatGPT-4o generates a heuristic decision tree for action selection.
  • VLM-DT (claude-3.5-sonnet): An equivalent baseline using Claude 3.5 Sonnet as the prompting engine.
@@
We use Proximal Policy Optimization (PPO) as the reinforcement learning algorithm to train our policy [24], with a learning rate of 0.0005 and a discount factor of 0.99. During testing, the policy is tested five times using different random seeds, each distinct from those used during training and evaluation. Results are shown in Fig. 3. Across all three environments, VLM-ICT consistently outperforms other uninterpretable models (MLP) built on the semantic feature space, and demonstrates performance close to the CNN+RL black-box benchmark representing the upper limit of task performance. These consistent results highlight the robustness and effectiveness of our proposed method.
Building on a semantically meaningful feature set, our ICTs contains nodes directly corresponding to human-understandable concepts. While we do not evaluate interpretability directly, this approach is consistent with prior work where performance serves as a proxy for meaningful representations [14, 36]. A previous human study [18] also supports the interpretability of such trees. We visualize the learned ICT in Appendix A.1, along with example cases demonstrating how the ICT yields plausible local explanations in example scenarios.
5 Limitations and future works
@@
Our framework is yet limited by the assumption that SAM performs reasonably well across environments, which may hinder generalizability in domains with poor segmentation. Future work could explore RL-based backpropagation to fine-tune VLMs, allowing them to iteratively improve outputs based on agent feedback [1]. This approach has the potential to enable VLMs to identify discrepancies between the underlying physics simulation and its visual representation, leading to improved feature extraction for downstream RL tasks.
6 Conclusion